OpenStudy (anonymous):

c=650n+600 repair c=450n+1800 replacement find the number of years for which the total cost of repair would equal the total cost of replacement.

hero (hero):

Set c = c and solve for n

OpenStudy (anonymous):

how? can you show me?

hero (hero):

c = 650 + 600 c = 450n + 1800 c = c 650n + 600 = 450n + 1800 200n = 1200 n = 1200/200 n = 6

hero (hero):

6 years

hero (hero):

This is about the easiest version of a systems of equations there is.

OpenStudy (anonymous):

okay i got it it's just that i got a lil confused i mean it's a word problem

hero (hero):

This is one of the easiest word problems

hero (hero):

associated with systems of equations

hero (hero):

I'm scared for you